Hook Industries' capital structure consists solely of debt and common equity. It can issue debt at Rd=11%, and its common stock currently pays a $2.00 dividend per share (Do=$2.00). The stock's price is currently $24.75, its dividend is expected to grow at a constant rate of 7% per year, its tax rate is 35%, and its WACC is 13.95%. What percentage of the company's capital structure consists of debt?
Answer #1

COst of equity = [ D1 / P0 ] +g

D1 = D0 (1+g)

= $ 2 ( 1 + 0.07 )

= $ 2 (1.07)

= 2.14

COst of equity = [ D1 / P0 ] +g

= [ 2.14 / 24.75 ] + 0.07

= 0.0865 + 0.07

= 0.1565

After Tax Cost of debt = rd * (1- T)

= 11% * [ 1 - 0.35]

= 11% * 0.65

= 7.15%

Let X be the weight of Debt in Capital STructure and "1-x" be the weight of Equity.

Source Weight Cost Wtd Cost
Debt X 7.15% 0.0715X
Equity 1 - X 15.65% 0.1565 - 0.1565X
WACC 0.1565 - 0.085X

Given WACC is 0.1395

Thus 0.1565 - 0.085X = 0.1395

= 0.017 = 0.085X

X = 0.017 / 0.085

= 0.20

weight of Debt is 20%